# Parc Científic UV, c/ Catedrático José Beltrán, 2, E-46980 Paterna (València), Spain

### v2

@@ 5 The complete mass matrix with right-handed neutrinos: Discussion and Outlook
Although deriving a generic parametrisation valid outside of the seesaw limit using this approach would definitely be very interesting, it is beyond the scope of this work. For other parametrisations in models with right-handed neutrinos, which are exact and valid outside of the seesaw limit, see Refs. Blennow:2011vn ; Donini:2012tt . In the next section and in App. A we discuss the seesaw limit, which of course should give rise to the standard CI at leading order.
5.1 The standard Casas-Ibarra parametrisation in the seesaw limit
@@
$W\;\simeq\;\begin{pmatrix}(\mathbb{I}_{n_{L}}-\tfrac{1}{2}\,\Theta\,\Theta^{\dagger})\,U&\Theta\\[4.0pt] -\Theta^{\dagger}U&\mathbb{I}_{n_{R}}-\tfrac{1}{2}\,\Theta^{\dagger}\Theta\,,\end{pmatrix}\;+\;\mathcal{O}(\Theta^{3})\,\qquad\Theta=m_{D}\,m_{R}^{-1}\in\mathbb{C}^{n_{L}\times n_{R}},$ (103)
@@
$m_{\text{light}}\simeq\,m_{D}m_{R}^{-1}m_{D}^{T}\in\mathbb{C}^{n_{L}\times n_{L}},\quad m_{\text{heavy}}\simeq m_{R}\in\mathbb{C}^{n_{R}\times n_{R}}\,.$ (104)
@@ 6 Conclusions
While this work was being finalised, Ref. Herrero-Brocal:2025zpb appeared on the arXiv, proposing a generalisation of the Casas–Ibarra parametrisation and demonstrating its equivalence to the master parametrisation of Refs. Cordero-Carrion:2018xre ; Cordero-Carrion:2019qtu . We emphasise that Ref. Herrero-Brocal:2025zpb pursues a distinct aim: it focuses on conditions that reduce the number of free parameters, whereas our emphasis is on generality and on applications to widely studied models. The new model we propose is, in fact, motivated by this broader perspective.
@@ Acknowledgements.
We are grateful for useful discussions with Pilar Hernández, Jacobo López-Pavón, Nuria Rius and Alejandro Ibarra. We also acknowledge discussing with Avelino Vicente at FLASY 2025 regarding our insight of a generalisation of the Casas–Ibarra parametrisation. SM would like to thank Federico Mescia for the kind hospitality at the Laboratori Nazionali di Frascati (LNF-INFN) during the early stages of this work. This work is partially supported by the Spanish Agencia Estatal de Investigación MICINN/AEI (10.13039/501100011033) grant PID2023-148162NB-C21 and the Severo Ochoa project MCIU/AEI CEX2023-001292-S. JHG is supported by the Consolidación Investigadora Grant CNS2022-135592, funded also by European Union NextGenerationEU / PRTR, and the Generalitat Valenciana Plan GenT Excellence Program (CIESGT2024/7). SM acknowledges financial support from the Consolidación Investigadora Grant CNS2022-136005. JR thanks the support by a grant Consolidar-2023-2027 from the Secretaría de Ciencia y Tecnología (SeCyT) de la Universidad Nacional de Córdoba (UNC) and also support from the Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET), Argentina. DV is supported by a McDonald Institute Theory Fellowship funded from the Canada First Research Excellence Fund through the Arthur B. McDonald Canadian Astroparticle Physics Research Institute, and a Subatomic Physics Discovery Grant (individual) from the Natural Sciences and Engineering Research Council of Canada. All Feynman diagrams were generated using the TikZ-Feynman package for  Ellis:2016jkw .

### v3

@@ V The complete mass matrix with right-handed neutrinos: Discussion and Outlook
Although deriving a generic parametrisation valid outside of the seesaw limit using this approach would definitely be very interesting, it is beyond the scope of this work. For other parametrisations in models with right-handed neutrinos, which are exact and valid outside of the seesaw limit, see Refs. [16, 38]. In the next section and in App. A we discuss the seesaw limit, which of course should give rise to the standard CI at leading order.
V.1 The standard Casas-Ibarra parametrisation in the seesaw limit
@@
$W\;\simeq\;\begin{pmatrix}(\mathbb{I}_{n_{L}}-\tfrac{1}{2}\Theta\Theta^{\dagger})\,U&\Theta\\ -\Theta^{\dagger}U&\mathbb{I}_{n_{R}}-\tfrac{1}{2}\,\Theta^{\dagger}\Theta\,\end{pmatrix}\;+\;\mathcal{O}(\Theta^{3})\,$,
$\Theta=m_{D}\,m_{R}^{-1}\in\mathbb{C}^{n_{L}\times n_{R}},$ (101)
@@
$m_{\text{light}}\simeq\,m_{D}m_{R}^{-1}m_{D}^{T}\in\mathbb{C}^{n_{L}\times n_{L}}\,$,
$m_{\text{heavy}}\simeq m_{R}\in\mathbb{C}^{n_{R}\times n_{R}}\,.$ (102)
@@ VI Conclusions
While this work was being finalised, Ref. [58] appeared on the arXiv, proposing a generalisation of the Casas–Ibarra parametrisation and demonstrating its equivalence to the master parametrisation of Refs. [31, 32]. We emphasise that Ref. [58] pursues a distinct aim: it focuses on conditions that reduce the number of free parameters, whereas our emphasis is on generality and on applications to widely studied models. The new model we propose is, in fact, motivated by this broader perspective.
@@ Acknowledgements.
We are grateful for useful discussions with Pilar Hernández, Jacobo López-Pavón, Nuria Rius and Alejandro Ibarra. We also acknowledge discussing with Avelino Vicente at FLASY 2025 regarding our insight of a generalisation of the Casas–Ibarra parametrisation. SM would like to thank Federico Mescia for the kind hospitality at the Laboratori Nazionali di Frascati (LNF-INFN) during the early stages of this work. This work is partially supported by the Spanish Agencia Estatal de Investigación MICINN/AEI (10.13039/501100011033) grant PID2023-148162NB-C21 and the Severo Ochoa project MCIU/AEI CEX2023-001292-S. JHG is supported by the Consolidación Investigadora Grant CNS2022-135592, funded also by European Union NextGenerationEU / PRTR, and the Generalitat Valenciana Plan GenT Excellence Program (CIESGT2024/7). SM acknowledges financial support from the Consolidación Investigadora Grant CNS2022-136005. JR thanks the support by a grant Consolidar-2023-2027 from the Secretaría de Ciencia y Tecnología (SeCyT) de la Universidad Nacional de Córdoba (UNC) and also support from the Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET), Argentina. DV is supported by a McDonald Institute Theory Fellowship funded from the Canada First Research Excellence Fund through the Arthur B. McDonald Canadian Astroparticle Physics Research Institute, and a Subatomic Physics Discovery Grant (individual) from the Natural Sciences and Engineering Research Council of Canada. All Feynman diagrams were generated using the TikZ-Feynman package for  [40].